# SPECIAL RELATIVITY

### v11

@@ 4.3 A Moving Emitter: Lorentz Transformations-Based Explanation
Another clear way to explain the aberration of light from a moving source is by using a Lorentz boost, which accounts for the uniform translational motion of the source in the laboratory frame. Maxwell’s equations are invariant under Lorentz transformations; however, the Lorentz boost introduces a time coordinate transformation of the form $t^{\prime}=t-xv/c^{2}$, which modifies the timing of events across space.
As a result, features such as the maximum of the electric field reach different positions in the lab frame at different times. For example, the field maximum at a position $x$ along the $x$-axis appears delayed compared to the maximum at $x=0$, with a time shift given by $\Delta t=xv/c^{2}$.
This spatially dependent time delay effectively causes a tilt in the phase front of the radiation. To first order in $v/c$, this corresponds to a rotation of the phase front by an angle approximately equal to $v/c$.
According to Maxwell’s electrodynamics, coherent radiation is always emitted perpendicular to the radiation phase front, as Maxwell’s equations exhibit no intrinsic anisotropy. Thus, when the uniform motion of the source is analyzed via Lorentz transformations, the aberration of light can be understood in the framework of relativistic kinematics. In this description, an additional phase chirp, $d\phi/dx=k_{x}=v\omega_{i}/c^{2}$, emerges, causing the array of identical elementary sources in the moving emitter to acquire different phases. Consequently, the plane wavefront undergoes a rotation under the Lorentz transformation, resulting in a radiated light beam propagating at an angle $v/c$. This yields the well-known phenomenon of aberration of light (Fig. 1, right).
We now consider the group velocity of the radiated beam. For a plane wave of the form $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$, the dispersion equation in Maxwell’s electrodynamics simplifies to $k_{z}^{2}+k_{x}^{2}-\omega^{2}/c^{2}=0$. Using the initial conditions and the Lorentz transformation, we obtain the following expression: $\omega=(\omega_{i}+vK_{\perp})$, $k_{z}=\omega_{i}/c$, $k_{x}=(v\omega_{i}/c^{2}+K_{\perp})$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated current density, $\omega_{i}$ is the frequency of the emitter radiation before the acceleration. Substituting these expressions into the dispersion equation, we verify that it is indeed satisfied, as expected.
@@
It is worth examining the significance of the fact that two different approaches—covariant and non-covariant—lead to the same group velocity. The key insight is that both methods accurately describe the same physical reality. Since the group velocity has an objective, convention-independent meaning, both approaches must necessarily yield the same physical results.
The distinction between absolute time coordinatization and Lorentz coordinatization is particularly intriguing. In Chapter 3, we explored how one can transform from absolute time coordinatization to Lorentz coordinatization. Specifically, manipulations involving the rule-clock structure in the lab frame can be understood as a simple change of the time variable, given by the transformation $t\to t+xv/c^{2}$. When combined with the Galilean transformation, this variable change ultimately leads to the Lorentz transformation in the context of absolute time coordinatization in an inertial frame. However, this change of variables has no intrinsic physical significance. Notably, there is a direct link between this time shift and the issue of clock synchronization. Importantly, the convention-independent results of calculations remain unchanged in the transformed variables. Consequently, there is no need to convert the results of electrodynamics problem-solving back into the original (3+1) variables.
From a pedagogical perspective, studying electrodynamics through variable transformations provides valuable insight. It is worth noting that the lack of a dynamical explanation for wavefront rotation in Lorentz coordinatization has been a point of concern for some physicists. However, as discussed in Chapter 3, a useful way to conceptualize wavefront rotation is to view it as a consequence of transitioning to a new time variable within the framework of Galilean (”single-frame”) electrodynamics.
@@ 4.4 Reflection from a Mirror Moving Parallel to Its Surface
Up to this point, the previous two sections have not introduced any fundamentally new results. As discussed earlier, when the sources are independent, adopting Lorentz coordinates poses no difficulty—each source can be assigned its own coordinate system, complete with an independent set of clocks. However, the situation becomes more complex when a secondary source interacts with light emitted by a primary source. In such cases, it becomes necessary to apply electrodynamics within an absolute time framework.
With the results from Chapter 3 now at our disposal, we are in a position to examine the intriguing phenomenon of radiation from moving sources.
@@
To examine this claim, we analyze the reasoning found in standard texts 121212 Many sources claim that no aberration occurs when light reflects off a mirror moving parallel to its surface. For example, Sommerfeld [23] states: ”For a mirror moving tangentially to its surface, the law of reflection that holds for a stationary mirror is preserved.”. Textbooks typically analyze reflection using two Lorentz reference frames. The fixed (lab) frame $S$ is at rest relative to the plane-wave emitter, while the moving frame $S^{\prime}$ travels at velocity $v$, in which the mirror is at rest. Both frames use a Cartesian coordinate system with the $x-y$ plane tangent to the mirror’s surface. The $x$-axis aligns with the mirror’s velocity $v$. For simplicity, we consider light incident from the $z$-direction in the lab frame. The incident wave is described by its four-dimensional wave vector, where the time-like component represents the angular frequency $\omega$, and the space-like components define the propagation direction. In the lab frame $(t,x,y,z)$, the wave vector has components $k_{1}=(\omega,0,0,-\omega/c)$, with the negative sign indicating propagation toward the mirror (see Fig. 3a). Our goal is to determine the wave vector of the reflected beam.
@@
Notably, in absolute time coordinatization, the wave equation explicitly depends on the velocity vector. Consequently, the solution naturally involves light beams of different frequencies, incorporating the Doppler effect. For a tangentially moving mirror with a finite aperture, angular frequency dispersion—an effect of first order in $v/c$—cannot be neglected.
@@
There is an intuitively plausible reason why textbook authors conclude that light reflected from a tangentially moving mirror experiences no aberration. A comparison between conventional aberration theory for light and sound theory can provide insight. In Section 4.15, we will consider a sound emitter at rest in the atmospheric frame and examine a mirror moving tangentially along its surface with velocity $v$. According to sound theory, the reflected beam’s energy transport does not deviate because the group velocity equals the phase velocity. When a plane wave falls the mirror normally, it produces an oblique sound beam. The (diagonal) wave equation for the moving mirror in the atmospheric frame is identical to that of an emitter at rest. Textbooks often draw important parallels between sound and light aberration in the first-order approximation  [^13] We should make one additional remark regarding two common errors when solving problems related to reflection from a mirror moving tangentially to its surface. These errors are not independent. Textbooks typically assume that the metric of the moving mirror is diagonal, similar to the metric of a ’plane wave’ emitter at rest, and that the group velocity equals the phase velocity. As a result, regardless of the mirror’s size, there is no deviation in the energy transport of the reflected light beam.. This, however, is a misconception. There is a fundamental difference between the propagation of light and sound. A proper treatment of light aberration must be based on the pseudo-Euclidean geometry of space-time.
@@ 4.5 Solving the Emitter-Mirror Problem in (3+1) Space-Time
Let us analyze the reflection of light from a mirror moving parallel to its surface, relying solely on relativistic kinematics. After applying a Galilean boost in the case of absolute time coordinatization, we employed the metric given by Eq. 11, which describes the electrodynamics of a moving light source as observed from an inertial frame.
From the structure of Eq. 11, it becomes evident that the electromagnetic field configuration — expressed in terms of the coordinates $x$ and $t$ of a detector at rest in the inertial frame—coincides with the field configuration at position $x-vt$ at time $t=0$.
Using a Fourier approach, we showed that the problem could be reduced to the case of reflection from a tangentially moving grating. We recall that with a plane wave $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ with the wavenumber vector $\vec{k}$ and the frequency $\omega$ dispersion equation in the absolute time coordinatization becomes: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$.
Within a single inertial frame, the equivalence between active and passive Galilean boosts implies that translating the system in one direction is indistinguishable from shifting the coordinate system in the opposite direction by the same amount. According to this equivalence, Maxwell’s equations remain valid in the comoving coordinate system. The wavenumber vector of the diffraced wave in the comoving coordinates is fixed by initial conditions. In fact, $k_{z}=\omega_{i}/c$, $k_{x}=K_{\perp}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated reflectance. Using the initial conditions and the Galilean transformation back to the lab frame, we obtain the following expression $\omega=\omega_{i}+K_{\perp}v$, $k_{z}=\omega_{i}/c$, $k_{x}=K_{\perp}$. Substituting there expressions into dispersion equation, we verify that it is satisfied, as expected.
We find the requirement that the wavenumber $K_{\perp}$ and the frequency change $\Delta\omega$ are related by $\Delta\omega=K_{\perp}v$. This phenomenon is interpreted as a Doppler effect, as the reflection occurs from a moving, rather than stationary, system. As we consider aberration effects only to first order in $v/c$, and under the assumption $cK_{\perp}/\omega_{i}\ll v/c$, no second-order terms appear in the expression for $\Delta\omega$.
@@
[…]
@@
A common misconception in the literature asserts that the direction of propagation in reflection from a tangentially moving mirror is not determined by the normal to the wavefront [^15] A widespread misunderstanding stems from assuming that the orientation of radiation wavefronts has objective significance. As Norton [24] notes: ”One might try to escape the problem by supposing that the direction of propagation is not always given by the normal to the wavefront. We might identify the direction of propagation with the direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation. Whatever may be the merits of such proposals, they are unavailable to some trying to implement a principle of relativity. If the direction of propagation of a plane wave is normal to the wavefronts in one inertial frame, then that must be true in all inertial frames.”. This misunderstanding arises from a failure to distinguish between convention-dependent and convention-invariant aspects of the theory. The direction of energy transport is an objective, convention-invariant quantity. In contrast, the phase front orientation—corresponding to an observer’s plane of simultaneity—lacks objective meaning, as no experimental method can precisely determine this orientation due to the finite speed of light.
@@ 4.6 A Moving Mirror: Lorentz Transformations-Based Explanation
To maintain a Lorentz coordinate system for a moving mirror in the lab frame, clock resynchronization is necessary. This entails a time shift of the form given in Eq. 13, namely $t\to t+xv/c^{2}$, which is valid to first order in $v/c$. This time shift corresponds to an effective rotation of the phase front of the incoming radiation by an angle of $v/c$. Within this framework, the mirror’s reflection is analyzed using Maxwell’s equations.
@@
It is also insightful to analyze the electrodynamics of an emitter at rest. A rotation of the $x$-axis without changing the orientation of the $t$-axis is described by a Galilean transformation $x^{\prime}=x-vt$, $t^{\prime}=t$. Conversely, a rotation of the $t$-axis without changing the orientation of the $x$-axis is described by the transformation $x^{\prime}=x$, $t^{\prime}=t+xv/c^{2}$. To first-order, this transformation mathematically is similar to the Galilean transformation $x^{\prime}=x+vt$, $t^{\prime}=t$.

### v12

@@ 4.3 A Moving Emitter: Lorentz Transformations-Based Explanation
An alternative and insightful approach to understanding the aberration of light from a moving source is through Lorentz coordinatization. This method provides a relativistically consistent explanation of the phenomenon within the framework of Maxwell’s equations.
This explanation utilizes a clock-re-synchronization procedure. The time coordinate $t_{L}$ in the lab frame, under Lorentz coordinatization, is obtained by introducing a time offset of the form given in Eq. (13).
Although Maxwell’s equations are invariant under Lorentz transformations, the Lorentz boost introduces a transformation of the time coordinate of the form $t^{\prime}=t-xv/c^{2}$. This spatially varying time shift introduces a delay in the emission of radiation across the source plane, effectively tilting the phase front of the emitted wave. To first order in $v/c$, this corresponds to a rotation of the phase front by an angle approximately equal to $v/c$.
This rotation manifests physically in the fact that features of the wave, such as the electric field maximum, reach different spatial positions at different times in the lab frame. For instance, the field maximum at position $x$ along the $x$-axis is delayed relative to the field maximum at $x=0$, with a time shift given by $\Delta t=xv/c^{2}$.
In accordance with Maxwell’s electrodynamics, coherent radiation is always emitted perpendicular to the phase front. Since Maxwell’s equations contain no intrinsic directional bias, the tilt in the wavefront—induced by the Lorentz time transformation—results in a corresponding shift in the direction of energy propagation. Within this relativistic description, the moving emitter exhibits a phase chirp across its aperture, given by $d\phi/dx=k_{x}=v\omega_{i}/c^{2}$, which means that the individual elementary sources now oscillate with position-dependent phase. As a result, the wavefront undergoes a rotation, and the emitted radiation propagates at an angle $\theta_{a}=v/c$, capturing the essence of the aberration effect (see Fig. 1, right).
We now examine the group velocity of the emitted beam. For a plane wave described by $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$, the dispersion relation in vacuum, derived from Maxwell’s equations, is $k_{z}^{2}+k_{x}^{2}-\omega^{2}/c^{2}=0$. In special relativity, the four quantities $(\omega,k_{x},k_{y},k_{z})$ form a four-vector and transform accordingly under Lorentz transformations. We will restrict our analysis to terms up to first order in $v/c$. Using the initial conditions and the Lorentz transformation, we obtain the following expression: $\omega=(\omega_{i}+vK_{\perp})$, $k_{z}=\omega_{i}/c$, $k_{x}=(v\omega_{i}/c^{2}+K_{\perp})$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated current density, $\omega_{i}$ is the frequency of the emitter radiation before the acceleration. Substituting these into the dispersion relation confirms that it remains satisfied, as expected.
@@
It is important to recognize the significance of the fact that both covariant (Lorentz-based) and non-covariant (absolute time-based) approaches yield the same group velocity. This agreement underscores the physical objectivity of the result: group velocity is a measurable, convention-independent quantity, and must therefore remain invariant regardless of the chosen coordinatization.
The distinction between absolute time coordinatization and Lorentz coordinatization, as discussed in Chapter 3, is particularly illuminating. For non-relativistic velocities, the transition between these two frameworks can be understood as a simple redefinition of the time variable: $t\to t+xv/c^{2}$.
When combined with the Galilean transformation, this leads naturally to the Lorentz transformation, even within the context of an absolute-time formalism. However, this change of variables has no intrinsic physical significance—it represents merely a reparametrization of time, not a real physical effect.
This coordinate transformation is closely tied to the issue of clock synchronization. Crucially, the physical content of electrodynamics remains unchanged under such transformations. As a result, once a solution is obtained in the transformed variables, there is no need to revert to the original (3+1) variables; the results remain valid and physically meaningful.
It is commonly assumed that the theory of relativity can be applied to physical processes without a detailed understanding of the clock synchronization procedure. Most textbooks suggest that an operational interpretation of Lorentz coordinatization within a single inertial frame is unnecessary.
Nevertheless, distinguishing between absolute-time and Lorentz coordinatization (from an operational standpoint) deepens conceptual understanding—even in the simple case of a single moving source. While some physicists have expressed concern about the lack of a ”dynamical” explanation for wavefront rotation in Lorentz coordinates, Chapter 3 offers a compelling perspective: the rotation arises naturally from the transformation of the time coordinate, which alters the simultaneity structure of spacetime events. Within the absolute-time framework, this provides an intuitive—though coordinate-dependent—explanation of the aberration phenomenon.
@@ 4.4 Reflection from a Mirror Moving Parallel to Its Surface
Until now, the previous two sections have not introduced any fundamentally new results. As previously discussed, when the sources are independent, adopting Lorentz coordinates presents no difficulty—each source can be assigned its own coordinate system, along with an independent set of clocks. The situation becomes more complex, however, when a secondary source interacts with light emitted by a primary source within a single inertial frame. This challenge in special relativity can be effectively addressed using an electrodynamics-based approach framed within an absolute time perspective using metric Eq. (11).
As noted earlier, the metric in Eq. (11) has not been previously reported in the literature. Existing studies generally assume—incorrectly—that both sources share the same Minkowski metric and are governed by Maxwell’s equations within a single inertial (laboratory) frame. This assumption is often treated as self-evident. In effect, the conventional approach relies on a hidden premise: that the principle of equivalence between active and passive boosts does not holds within a single inertial frame. Our formulation challenges this assumption.
The agreement between our theoretical framework and experimental observations lends strong support to our revised formulation of special relativity, which incorporates the metric of Eq. (11) (see Section 4.14 for further discussion).
With the results of Chapter 3 now established, we are prepared to examine the intriguing phenomenon of radiation from moving sources.
@@
To examine this claim, we analyze the reasoning found in standard texts. 121212 Many sources claim that no aberration occurs when light reflects off a mirror moving parallel to its surface. For example, Sommerfeld [23] states: ”For a mirror moving tangentially to its surface, the law of reflection that holds for a stationary mirror is preserved.” Textbooks typically analyze reflection using two Lorentz reference frames. The fixed (lab) frame $S$ is at rest relative to the plane-wave emitter, while the moving frame $S^{\prime}$ travels at velocity $v$, in which the mirror is at rest. Both frames use a Cartesian coordinate system with the xyx-y plane tangent to the mirror’s surface. The $x$-axis aligns with the mirror’s velocity $v$. For simplicity, we consider light incident from the $z$-direction in the lab frame. The incident wave is described by its four-dimensional wave vector, where the time-like component represents the angular frequency $\omega$, and the space-like components define the propagation direction. In the lab frame $(t,x,y,z)$, the wave vector has components $k_{1}=(\omega,0,0,-\omega/c)$, with the negative sign indicating propagation toward the mirror (see Fig. 3a). Our goal is to determine the wave vector of the reflected beam.
@@
Notably, in absolute time coordinatization, the wave equation explicitly depends on the velocity vector. Consequently, the solution naturally involves light beams of different frequencies, incorporating the Doppler effect. For a tangentially moving mirror with a finite aperture, angular frequency dispersion—an effect of first order in v/cv/c—cannot be neglected.
@@
There is an intuitively plausible reason why textbook authors conclude that light reflected from a tangentially moving mirror experiences no aberration. A comparison between conventional aberration theory for light and sound theory can provide insight. In Section 4.15, we will consider a sound emitter at rest in the atmospheric frame and examine a mirror moving tangentially along its surface with velocity $v$. According to sound theory, the reflected beam’s energy transport does not deviate because the group velocity equals the phase velocity. When a plane wave falls the mirror normally, it produces an oblique sound beam. The (diagonal) wave equation for the moving mirror in the atmospheric frame is identical to that of an emitter at rest. Textbooks often draw important parallels between sound and light aberration in the first-order approximation.  [^13] We should make one additional remark regarding two common errors when solving problems related to reflection from a mirror moving tangentially to its surface. These errors are not independent. Textbooks typically assume that the metric of the moving mirror is diagonal, similar to the metric of a ’plane wave’ emitter at rest, and that the group velocity equals the phase velocity. As a result, regardless of the mirror’s size, there is no deviation in the energy transport of the reflected light beam. This, however, is a misconception. There is a fundamental difference between the propagation of light and sound. A proper treatment of light aberration must be based on the pseudo-Euclidean geometry of space-time.
@@ 4.5 Solving the Emitter-Mirror Problem in (3+1) Space-Time
Let us analyze the reflection of light from a mirror moving tangentially (i.e., parallel to its surface), relying solely on relativistic kinematics. To describe this scenario, we adopt the framework of absolute time coordinatization and apply a Galilean boost. In this context, we utilize the metric given by Eq. 11, which characterizes the electrodynamics of a moving light source as observed from an inertial frame.
From the structure of Eq. 11, it becomes evident that the electromagnetic field configuration—expressed in terms of the coordinates $x$ and $t$ of a detector at rest in the inertial frame—coincides with the field configuration at position xvtx-vt at time $t=0$.
Using a Fourier analysis, we show that the problem can be reduced to that of reflection from a tangentially moving diffraction grating. Consider a plane wave of the form $\exp(i\vec{k}\cdot\vec{r}-i\omega t)$ where $\vec{k}$ is the wavevector and $\omega$ the angular frequency. In the absolute time coordinatization, the dispersion relation takes the form: $(1-v^{2}/c^{2})k_{x}^{2}+2vk_{x}\omega/c^{2}+k_{z}^{2}-\omega^{2}/c^{2}=0$.
Within a single inertial frame, the equivalence between active and passive Galilean boosts implies that translating the system in one direction is indistinguishable from shifting the coordinate system in the opposite direction by the same amount. This equivalence ensures that Maxwell’s equations retain their form in the comoving coordinate system.
In the comoving frame, the wavevector of the diffracted wave is determined by the initial conditions. Specifically, we have: $k_{z}^{2}=\omega_{i}^{2}/c^{2}-k_{x}^{2}$, $k_{x}=K_{\perp}$, where $K_{\perp}$ denotes the spatial frequency of the sinusoidally space-modulated reflectance. Transforming back to the lab frame using a Galilean transformation, we find that the frequency shifts according to: $\omega=\omega_{i}+K_{\perp}v$, while the components of the wavevector remain unchanged.
Substituting these expressions into the dispersion relation yields: $(1-v^{2}/c^{2})K_{\perp}^{2}+2vK_{\perp}(\omega_{i}+K_{\perp}v)/c^{2}+\omega_{i}^{2}/c^{2}-K_{\perp}^{2}-(\omega_{i}+K_{\perp}v)^{2}/c^{2}=0$, which is readily verified to be satisfied, as expected.
This analysis leads to the conclusion that the frequency shift $\Delta\omega$ and the spatial frequency $K_{\perp}$ are related by: $\Delta\omega=K_{\perp}v$. This relation is recognized as a manifestation of the Doppler effect, arising from reflection off a moving (as opposed to stationary) surface.
@@
One of the consequences of the Doppler effect is the angular frequency dispersion of light waves reflected from a moving mirror with a finite aperture. If $\vec{n}=\vec{k}/|\vec{k}|$ denotes a unit vector in the direction of the wave normal, and $\vec{v}$ is the mirror velocity vector relative to the lab frame, we get the equation $\omega_{s}=\omega_{i}(1+\vec{n}\cdot\vec{v}/c)=\omega_{i}+(\omega_{i}v/c)\cos\theta$. The Doppler effect is responsible for angular frequency dispersion to the first order of $v/c$ even when $\vec{n}\cdot\vec{v}=0$ (i.e when $\cos\theta=0$). In fact, $\d\omega_{s}/d\theta=-(\omega_{i}v/c)\sin\theta=-\omega_{i}v/c$ at $\theta=\pi/2$. We can rewrite this equation in a different way. The differential of the scattered angle is given by $d\theta=-dk_{x}/k_{i}$. With the help of this relation and account for that $k_{i}=\omega_{i}/c$ we have $d\omega_{s}/dk_{x}=v$.
@@
A common misconception in the literature asserts that the direction of propagation in reflection from a tangentially moving mirror is not determined by the normal to the wavefront. [^15] A widespread misunderstanding stems from assuming that the orientation of radiation wavefronts has objective significance. As Norton [24] notes: ”One might try to escape the problem by supposing that the direction of propagation is not always given by the normal to the wavefront. We might identify the direction of propagation with the direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation. Whatever may be the merits of such proposals, they are unavailable to some trying to implement a principle of relativity. If the direction of propagation of a plane wave is normal to the wavefronts in one inertial frame, then that must be true in all inertial frames.” This misunderstanding arises from a failure to distinguish between convention-dependent and convention-invariant aspects of the theory. The direction of energy transport is an objective, convention-invariant quantity. In contrast, the phase front orientation—corresponding to an observer’s plane of simultaneity—lacks objective meaning, as no experimental method can precisely determine this orientation due to the finite speed of light.
@@ 4.6 A Moving Mirror: Lorentz Transformations-Based Explanation
Another clear way to explain the aberration of light from a moving mirror is by using a Lorentz boost, which accounts for the uniform translational motion of the mirror in the laboratory frame. Maxwell’s equations are invariant under Lorentz transformations; however, the Lorentz boost introduces a space-time coordinates transformation of the form Eq. (13). To minimize the mathematical complexity of the discussion, we will restrict our analysis to terms up to first order in $v/c$. This entails only a time shift of the form $t\to t+xv/c^{2}$. This time shift corresponds to an effective rotation of the phase front of the incoming radiation by an angle of $v/c$. Within this framework, the mirror’s reflection is analyzed using Maxwell’s equations.
@@
It is also insightful to analyze the electrodynamics of an emitter at rest in the case of Lorentz coordinatization. A rotation of the $x$-axis without changing the orientation of the $t$-axis is described by a Galilean transformation $x^{\prime}=x-vt$, $t^{\prime}=t$. Conversely, a rotation of the $t$-axis without changing the orientation of the $x$-axis is described by the transformation $x^{\prime}=x$, $t^{\prime}=t+xv/c^{2}$. To first-order, this transformation mathematically is similar to the Galilean transformation $x^{\prime}=x+vt$, $t^{\prime}=t$.